OpenStudy (anonymous):

If f(x) = x^2 + 3x - 4, then f(x+1) (a) x^2 + 3x - 3 (b) x(x+5) (c) x^2 + 5x + -3 (d) x^2 + 3x - 1 (e) 3x + 1

OpenStudy (anonymous):

replace x with x+1

Parth (parthkohli):

What you must find is \((x + 1)^2 + 3(x + 1) - 4\)

OpenStudy (anonymous):

i got x^2 + 5x. so that's (b) right?

hero (hero):

How come you can't just add 1 to x^2 + 3x - 4 ?

Parth (parthkohli):

Yes, it's correct. @lawls

Parth (parthkohli):

@Hero Nope. You have to review functions again.

hero (hero):

lol, I responded too quickly

Parth (parthkohli):

For example:\[f(x) = x + 5 \implies f(\theta) = \theta + 5\]Just substitute and play!

hero (hero):

My bad. I made a mistake.

OpenStudy (anonymous):

thank you :)

hero (hero):

I know what f(x+1) is. It was an honest mistake.

hero (hero):

I was thinking of f(x) + 1

Parth (parthkohli):

You're welcome @lawls :)

Parth (parthkohli):

@Hero Not to worry too much about it. lol I know you are a genius =)